User name jpchen
Log entry time 16:03:37 on February 6,2002
Entry number 81992
keyword=day shift summary
Day shift, 2/6/2002
Shift Workers: J. P. Chen, L. Pentchev, V. Koubarovski.
8:00 Shift started with data taking in Kin 5E(e).
Previous shift tried higher current (>30 uA), settled at 20 uA.
8:09 Run 2143, production.
8:40 Beam off, MCC called, one RF zone dropped, will be down for 10-15 minutes.
10:00 Beam back, resume data taking.
10:00-15:45 , Runs 2144-2164 data taking for 5E(e).
13:42 run 2157, TS2 not responding. kcoda and restarted.
14:35 run2160, TS2 not responding again. kcoda and restarted.
15:20 ROC14 crashed. kcoda, reset and restarted.
15:45 Completed 5E(e), switched to 5E(f), Run 2165.
16:00 transfer to next shift.